Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orlati andrea at Sat Feb 16 07:34:47 MET 2002) Minor failures
strong interference saturated video converter 1 and 2. scan between 00:01:30 ut day 47 and 02:06 ut were lost or partially lost due to system failure.

Torun (Kaz Borkowski at Mon Feb 18 09:48:05 MET 2002) Minor failures
Late start, at 22:45, due to problems with IF Distributors. A few scans,23:56 to 23:59 (day 46) and 1:09 to 1:15 (day 47), lost due to communication error (first case) and early tape end (a later consequence of the same).

Jodrell_Bank (Alastair Gunn at Mon Feb 18 11:07:23 MET 2002) Minor failures
Bad count readback in VC04, therefore Tsys corrupted. Voltage still looks okay on VC04 so data should be good. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 18 11:07:48 MET 2002) Failure
Cambridge did not observe due to serious fault with recorder.

Effelsberg (A. Kraus at Mon Feb 18 11:50:04 MET 2002) Success

Onsala (John Conway at Tue Feb 19 08:37:30 MET 2002) Success
No known problems

Westerbork (Tony Foley at Tue Feb 19 14:36:44 MET 2002) Minor failures
No known problems